OpenStudy (lukecrayonz):

Explain your process for writing the standard form equation for an ellipse.

OpenStudy (anonymous):

From what? Given information or general form?

OpenStudy (lukecrayonz):

I have no idea, that's all that it says :(

OpenStudy (lukecrayonz):

That's why I'm here haha

OpenStudy (anonymous):

Well, have you learned the form ax^2 + by^2 + cx + dy + f = 1 yet?

OpenStudy (lukecrayonz):

Yes

OpenStudy (anonymous):

I'll just explain that then.

OpenStudy (lukecrayonz):

well it's e, but you know that :P

OpenStudy (lukecrayonz):

Wait isnt that a hyperbola?:O not an ellipse?

OpenStudy (anonymous):

An ellipse is added in between.

OpenStudy (anonymous):

First predict what type of conic it is. If the squared terms are both positive, but different coefficients, then it's an ellipse. Let's use the example: 4x^2 + 9y^2 - 32x + 36y + 64 = 0 4x^2 + 9y^2 - 32x + 36y = -64 Then, rewrite the conic so that the x and the y together. 4x^2 - 32x + 9y^2 + 36y = -64 Factor out a 4 from the x terms. 4(x^2 - 8x) + 9y^2 + 36y = -64 Factor out a 9 from the y terms. 4(x^2 - 8x) + 9(y^2 + 4y) = -64 Complete the square for both the x and y terms and balance the equation by adding to both sides. 4(x^2 - 8x + 16) + 9(y^2 + 4y + 4) = -64 + 64 + 36 Add like terms. 4(x^2 - 8x + 16) + 9(y^2 + 4y + 4) = 36 Factor the perfect squares. 4(x - 4)^2 + 9(y + 2)^2 = 36 Divide both sides by 36. (x - 4)^2/9 + (y + 2)^2/4 = 1

OpenStudy (lukecrayonz):

Thank you so much!

OpenStudy (anonymous):

You're welcome.